OpenStudy (anonymous):

Help explain how to solve simultaneously the below equations: 10 = v cos theta + 5.638 4 = v sin theta - 2.05 I need to figure out angle theta and v?

OpenStudy (anonymous):

I need to know what to solve for in the 1st equation? I know it seems easy, I'm confusing myself a bit too much. Help!

OpenStudy (anonymous):

First you seperate the vcos thete & vsin theta & take their squre. After squaring add them to find v & the put the value of v in any equation you can get theta...

OpenStudy (anonymous):

A fellow user helped me with this: v cos theta = 4.362 and v sin theta = 6.05 I divide (v sin theta/v cos theta) = (6.05/4.362) v tan theta = 1.39 inverse tan of 1.39 = 54.21 degrees I substituted that into 1st equations = v (cos 54.21) = 4.362 v = 7.46 m/s

OpenStudy (anonymous):

\[10=v \cos \Theta+5.638\] \[\rightarrow v \cos \Theta =10-5.638\] \[\rightarrow v ^{2} \cos ^{2} \Theta=(4.362)^{2}\] \[\rightarrow v ^{2} \cos ^{2} \theta=19.027..................1 no.\] and then \[4=v \sin \Theta-2.05\] \[\rightarrow v ^{2} \sin ^{2} \Theta=(4+2.05)^{2}\] \[\rightarrow v ^{2} \sin ^{2} \Theta=36.603..................2 no\] then add 1no. and 2.no \[v^{2}\sin^{2}\Theta+v^{2}\cos^{2}\theta=19.027+36.603\] \[v^{2}=55.63\] \[v=\sqrt{55.63}\] \[v=7.459...(Answer)\]

OpenStudy (anonymous):

@biswajit_paul where is sin2theta and cos2theta go when we added them?

OpenStudy (anonymous):

is it a trig rule?

OpenStudy (anonymous):

@Boushra It comes from euclidean geometry. \[\sin^{2}\theta + \cos^{2}\theta=1\]

OpenStudy (anonymous):

It is trigonometry

OpenStudy (anonymous):

super! I will keep that in mind :) thank you so much!